### v2

@@ 3 Main Approximation Theorem
The quality of the approximation will be determined by the extent to which these maps commute. The neural operator $\mathcal{G}_{\theta}$ will approximate the ‘true map’ $\mathcal{G}^{\dagger}_{\text{ext}}$. Recall that we denote encoding (by sensor evaluations) by $\mathcal{E}$, decoding by $\mathcal{D}$, Reconstructor map by $\mathcal{R}$, the projector map by $\mathcal{P}$, and the approximator map between finite dimensional spaces as $\mathcal{A}$. Referring to Fig. 1, notice that our neural operator will have the form $\mathcal{G}_{\theta}=\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}$. Our goal is to find an upper bound on the error incurred in approximating $\mathcal{G}^{\dagger}_{\text{ext}}$ by $\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}:=\mathcal{G}_{\theta}$ over the space of shifted N-t-D operators corresponding to admissible conductivities. More particularly we will prove the following theorem.
Theorem 1.
@@
$\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert\mathcal{G}^{\dagger}_{\text{ext}}(\tilde{\Lambda}_{\gamma})-\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}(\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}\leq\epsilon$ (14)
Proof.
@@
$+\underbrace{\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}-\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}}_{E_{\mathcal{R}}})\bigg{)}$ (15)
where $E_{\mathcal{A}},E_{\mathcal{E}}$ and $E_{\mathcal{R}}$ denote the approximation error, encoding error, and the reconstruction error respectively. In the next three subsections we will analyze each of these error terms separately. In particular, we will show that for any given $\epsilon$ the network parameters can be chosen to ensure that each of the error terms separately is of the order $\epsilon/3$. Here and below, the symbol $\lesssim$ means that the inequality holds upto some constant factor, and the notation $\lVert\cdot\rVert_{L^{2}}$ is a shorthand for $\lVert\cdot\rVert_{L^{2}(\Omega)}$. Furthermore, the fact that $\Omega$ is a smooth bounded domain allows us to translate upper bounds in $L^{\infty}(\Omega)$ easily into upper bounds with respect to $L^{2}(\Omega)$ norm, we will use this fact freely wherever needed.
@@ 3.1 Approximator error
First we estimate the approximation error, $E_{\mathcal{A}}$. Recall that $\tilde{\Lambda}_{\gamma}$ is a bounded map for admissible conductivities, [26, Theorem A.3]. Thus
$E_{\mathcal{A}}:$ $=\lvert\lvert(\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}\lesssim\mathrm{Lip}(\mathcal{R})\lvert\lvert\mathcal{A}-\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\rvert\rvert_{l^{2}}\mathrm{Lip}(\mathcal{E})$
Here $\mathrm{Lip(\mathcal{R})}$ and $\mathrm{Lip(\mathcal{E})}$ are the Lipschitz constants for the (Lipschitz) maps $\mathcal{R}$ and $\mathcal{E}$. As the map $\mathcal{R}$ is similar to the one defined in [43], the Lipschitz constant of $\mathcal{R}$ follows from [43, Lemma 3.2]. Now if we consider the encoding map $\mathcal{E}(\tilde{\Lambda}_{\gamma}):=\tilde{R}^{M}_{\gamma}$, then we can show that this is Lipschitz, too. This follows from the discussion in [44, Sections 2 and 4] from which we first get that the map $\gamma\mapsto\tilde{R}^{M}_{\gamma}$ is Lipschitz and subsequently we apply the estimate ([27, Theorem 2.3]), $\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}\leq C\lvert\lvert\tilde{\Lambda}_{\gamma_{1}}-\tilde{\Lambda}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}$. Now the map, $\mathcal{G}^{\dagger}_{\text{ext}}$ is a Lipschitz continuous map for piecewise analytic conductivities, [27, Theorem 2.3]. Furthermore, the map $\mathcal{P}$ is Lipschitz by similar argument as in [43, Lemma 3.2, 3.3]. $\mathcal{D}$ is also Lipschitz from the following estimate and the fact that the projection operator $\tilde{P}^{E}_{M}$ is bounded,
@@
Thus the composition $\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}:=G:\mathbb{R}^{M}\to\mathbb{R}^{P}$ is also Lipschitz. The upper bound on the aproximation error can be evaluated by finding an upper bound for the error incurred in approximating the map $G:\mathbb{R}^{M}\to\mathbb{R}^{P};{}x\mapsto(G_{1}(x),\cdots,G_{P}(x))$ by a neural network $\mathcal{A}$. Clearly, this can be done by approximating each of the maps $G_{j}:\mathbb{R}^{M}\to\mathbb{R},{}j\in\{1,\cdots,P\}$ by a neural network $\mathcal{A}_{j}:\mathbb{R}^{M}\to\mathbb{R}$ and then combining all the individual approximations into a single neural network $\mathcal{A}$ wherein
@@
Then it follows from [73, Theorem 1] (see also discussion following [43, Theorem 3.17])that for any given $\epsilon/3:=\tilde{\epsilon}>0$, there exists a neural network $\mathcal{A}_{j}$ with
@@
$\lVert G_{j}(x)-\mathcal{A}_{j}(x)\rVert_{L^{\infty}([0,1])^{M}}\leq\tilde{\epsilon}.$ (16)
Thus there exists a neural network $\mathcal{A}$ whose size is of the order $\mathcal{O}(P\tilde{\epsilon}^{-M})$ that can approximate a Lipschitz continuous mapping $G:\mathbb{R}^{M}\to\mathbb{R}^{P}$ within a given $\epsilon/3=\tilde{\epsilon}$ error, see also [43, section 3.6.1].
@@ 3.2 Encoding error
$\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}=\lvert\lvert\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}((\mathcal{D}\circ\mathcal{E}-\mathbb{I}_{\mathbb{H}_{r}})\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}$ (17)
where $\mathbb{I}_{\mathbb{H}_{r}}$ is the identity map on $\mathbb{H}_{r}$. Observe that $(\mathcal{D}\circ\mathcal{E}-\mathbb{I}_{\mathbb{H}_{r}})\tilde{\Lambda}_{\gamma}=\tilde{\Lambda}^{M}_{\gamma}-\tilde{\Lambda}_{\gamma}$. From [20, Corollory 6.2], we know that by suitably shrinking the size of electrodes as their number $M$ grows we get
Lemma 1.
@@
$\lvert\lvert\tilde{\Lambda}_{\gamma}-\tilde{\Lambda}^{M}_{\gamma}\rvert\rvert_{\mathcal{L}(L^{2}_{\diamond}\to L^{2}_{\diamond})}:=\delta_{M}\to 0\quad\text{as}\quad M\to\infty.$ (18)
As a simple consequence to the lemma above, and similar to the calculations in [1, Lemma 5] one can establish the relation between operator norm $\mathcal{L}(L^{2}_{\diamond}\to L^{2}_{\diamond})$ and the information-theoretically relevant $\mathbb{H}_{r}$-norm to get $\lvert\lvert\tilde{\Lambda}_{\gamma}-\tilde{\Lambda}^{M}_{\gamma}\rvert\rvert_{\mathbb{H}_{r}}\to 0$ as $M\to\infty$. As is shown in Lemma 3 below, $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$. Consider the finite rank projection operators, $U_{M}:\mathbb{H}_{r}\to\mathbb{H}_{r}$ where $U_{M}:=\mathcal{D}_{M}\circ\mathcal{E}_{M}$ where $\mathcal{E}_{M}$ and $\mathcal{D}_{M}$ are as defined in section 2. From Lemma 1, we have that for any $\tilde{\Lambda}_{\Gamma}\in D_{\Lambda}$:
$\lim\limits_{M\to\infty}\sup\limits_{\tilde{\Lambda}_{\gamma}}\lvert\lvert\Lambda_{\gamma}-U_{M}(\tilde{\Lambda}_{\gamma})\rvert\rvert_{\mathbb{H}_{r}}=0.$ (19)
@@
$\forall z_{1},z_{2}\in Z,\quad\lvert\lvert\mathcal{G}^{\dagger}_{\text{ext}}(z_{1})-\mathcal{G}^{\dagger}_{\text{ext}}(z_{2})\rvert\rvert_{L^{\infty}}\leq\omega(\lvert\lvert z_{1}-z_{2}\rvert\rvert_{\mathbb{H}_{r}})$ (20)
@@
$\sup\limits_{z\in D_{\Lambda}}\omega(\lvert\lvert U_{M}(z)-z\rvert\rvert_{\mathbb{H}_{r}})\leq\epsilon/3.$ (21)
@@
$\sup\limits_{\tilde{\Lambda}_{\gamma}}\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}\lesssim\epsilon/3.$ (22)
@@ 3.3 Reconstruction error
Note that the output of the DeepONet lies in the space of $L^{2}$ functions. For any $\tilde{\Lambda}_{\gamma}\in D_{\Lambda}$, the third term in (3) can be expressed as $E_{\mathcal{R}}={\lvert\lvert(\mathcal{R}\circ\mathcal{P}-\mathbb{I}_{\Omega})(\mathcal{G}^{\dagger}_{\text{ext}}\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}}$, where $\mathbb{I}_{\Omega}$ represents the identity map on the space of $L^{2}(\Omega)$ functions. Denote $g:=\mathcal{G}^{\dagger}_{\text{ext}}(\tilde{\Lambda}_{\gamma})=\mathcal{G}^{\dagger}(\tilde{\Lambda}_{\gamma})=\Phi^{-1}\circ\gamma$, where $\Phi$ is the link function as described in section 2. Since $\gamma$ is a piecewise analytic function, hence it belongs to some $H^{s}(\Omega)$ for $20<s<1/$ 2. It is easy to see that,
@@
where $\tau_{0}^{\perp}:=\tau_{0}-\Sigma_{k=1}^{P}\langle\tau_{0},\tau_{k}^{*}\rangle\tau_{k}$ is the projection of $\tau_{0}$ onto the orthogonal complement of $\mathrm{span}(\tau_{1},\cdots,\tau_{p})$. Following arguments from [43, Lemma 3.2,3.3, 3.4], we have the following Lemma.
Lemma 2.

### v3

@@ 3 Main Approximation Theorem
The quality of the approximation will be determined by the extent to which the maps in Fig. 1 commute. The neural operator $\mathcal{G}_{\theta}$ will approximate the ‘true map’ $\mathcal{G}^{\dagger}_{\text{ext}}$. Recall that we denote encoding (by sensor evaluations) by $\mathcal{E}$, decoding by $\mathcal{D}$, Reconstructor map by $\mathcal{R}$, the projector map by $\mathcal{P}$, and the approximator map between finite dimensional spaces as $\mathcal{A}$. Referring to Fig. 1, notice that our neural operator will have the form $\mathcal{G}_{\theta}=\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}$. Our goal is to find an upper bound on the error incurred in approximating $\mathcal{G}^{\dagger}_{\text{ext}}$ by $\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}:=\mathcal{G}_{\theta}$ over the space of shifted N-t-D operators corresponding to admissible conductivities. More particularly we will prove the following theorem.
Theorem 3.1.
@@
(14) $\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert\mathcal{G}^{\dagger}_{\text{ext}}(\tilde{\Lambda}_{\gamma})-\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}(\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}\leq\epsilon$
Proof 3.2.
@@
(15) $+\underbrace{\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}-\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}}_{E_{\mathcal{R}}})\bigg{)}$
where $E_{\mathcal{A}},E_{\mathcal{E}}$ and $E_{\mathcal{R}}$ denote the approximator error, encoder error, and the reconstructor error respectively. In the next three subsections we will analyze each of these error terms separately. In particular, we will show that for any given $\epsilon$ the network parameters can be chosen to ensure that each of the error terms separately is of the order $\epsilon/3$. Here and below, the symbol $\lesssim$ means that the inequality holds upto some constant factor, and the notation $\lVert\cdot\rVert_{L^{2}}$ is a shorthand for $\lVert\cdot\rVert_{L^{2}(\Omega)}$. Furthermore, the fact that $\Omega$ is a smooth bounded domain allows us to translate upper bounds in $L^{\infty}(\Omega)$ easily into upper bounds with respect to $L^{2}(\Omega)$ norm, we will use this fact freely wherever needed.
@@ 3.1 Approximator error
First we estimate the approximation error, $E_{\mathcal{A}}$. To that end, we examine the continuity of various maps appearing in the error terms above. Recall that $\tilde{\Lambda}_{\gamma}$ is a bounded map for admissible conductivities, [27, Theorem A.3]. As the map $\mathcal{R}$ is similar to the one defined in [44], the Lipschitz constant of $\mathcal{R}$ follows from [44, Lemma 3.2]. Consider the encoding map $\mathcal{E}(\tilde{\Lambda}_{\gamma}):=\tilde{R}^{M}_{\gamma}$, then we can show that this is Lipschitz, too. This follows from the discussion in [45, Sections 2 and 4] from which we first get that the map $\gamma\mapsto\tilde{R}^{M}_{\gamma}$ is Lipschitz and subsequently we apply the estimate ([28, Theorem 2.3]), $\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}\leq C\lvert\lvert\tilde{\Lambda}_{\gamma_{1}}-\tilde{\Lambda}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}$. As $D_{\Lambda}$ is a compact set from Lemma 7.1, then $\mathcal{E}(D_{\Lambda})$ is also compact. Now the map, $\mathcal{G}^{\dagger}_{\text{ext}}$ is a Lipschitz continuous map for piecewise analytic conductivities, [28, Theorem 2.3]. Furthermore, the map $\mathcal{P}$ is Lipschitz by similar argument as in [44, Lemma 3.2, 3.3]. $\mathcal{D}$ is also Lipschitz from the following estimate and the fact that the auxiliary projection operator, $\tilde{P}^{E}_{M}$ is bounded, see e.g. [5].
@@
Thus the composition $\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}:=G:\mathbb{R}^{M\times M}\to\mathbb{R}^{P}$ is also Lipschitz. Here and below, let $\mathrm{Lip(\mathcal{R})}$ and $\mathrm{Lip(\mathcal{E})}$ denote the Lipschitz constants for the (Lipschitz) maps $\mathcal{R}$ and $\mathcal{E}$. Thus
$E_{\mathcal{A}}$ $\leq\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert(\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}-\mathcal{R}\circ G\circ\mathcal{E})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}$
$\leq\operatorname{Lip}(\mathcal{R})\cdot\sup_{X\in\mathcal{E}(D_{\Lambda})}\left\|\mathcal{A}(X)-G(X)\right\|,$
The upper bound on the approximation error can be evaluated by finding an upper bound for the error incurred in approximating the map $G:\mathbb{R}^{M\times M}\to\mathbb{R}^{P};{}X\mapsto(G_{1}(X),\cdots,G_{P}(X))$ by a neural network $\mathcal{A}$. This can be done using the results proved by Yarotsky in [75]. To make this precise, we begin by approximating each of the maps $G_{j}:\mathbb{R}^{M\times M}\to\mathbb{R},{}j\in\{1,\cdots,P\}$ by a neural network $\mathcal{A}_{j}:\mathbb{R}^{M\times M}\to\mathbb{R}$ and then combining all the individual approximations into a single neural network $\mathcal{A}$ wherein
@@
Then it follows from [75, Theorem 1] (see also discussion following [44, Theorem 3.17])that for any given $\epsilon/3:=\tilde{\epsilon}>0$, there exists a neural network $\mathcal{A}_{j}$ with
@@
(16) $\sup\limits_{X\in\mathcal{E}(D_{\Lambda})}\lVert G_{j}(X)-\mathcal{A}_{j}(X)\rVert\leq\tilde{\epsilon}.$
Thus there exists a neural network $\mathcal{A}$ whose size is of the order $\mathcal{O}(P\tilde{\epsilon}^{-M})$ that can approximate a Lipschitz continuous mapping $G:\mathbb{R}^{M\times M}\to\mathbb{R}^{P}$ within a given $\epsilon/3=\tilde{\epsilon}$ error, see also [44, section 3.6.1].
@@ 3.2 Encoding error
(17) $\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}=\lvert\lvert\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}((\mathcal{D}\circ\mathcal{E}-\mathbb{I}_{\mathbb{H}_{r}})\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}$
where $\mathbb{I}_{\mathbb{H}_{r}}$ is the identity map on $\mathbb{H}_{r}$. Observe that $(\mathcal{D}\circ\mathcal{E}-\mathbb{I}_{\mathbb{H}_{r}})\tilde{\Lambda}_{\gamma}=\tilde{\Lambda}^{M}_{\gamma}-\tilde{\Lambda}_{\gamma}$. From [21, Corollory 6.2], we know that by suitably shrinking the size of electrodes as their number $M$ grows we get
Lemma 3.3.
@@
(18) $\lvert\lvert\tilde{\Lambda}_{\gamma}-\tilde{\Lambda}^{M}_{\gamma}\rvert\rvert_{\mathcal{L}(L^{2}_{\diamond}\to L^{2}_{\diamond})}:=\delta_{M}\to 0\quad\text{as}\quad M\to\infty.$
As a simple consequence to the lemma above, and similar to the calculations in [1, Lemma 5] one can establish the relation between operator norm $\mathcal{L}(L^{2}_{\diamond}\to L^{2}_{\diamond})$ and the information-theoretically relevant $\mathbb{H}_{r}$-norm to get $\lvert\lvert\tilde{\Lambda}_{\gamma}-\tilde{\Lambda}^{M}_{\gamma}\rvert\rvert_{\mathbb{H}_{r}}\to 0$ as $M\to\infty$. As is shown in Lemma 7.1 below, $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$. Consider the finite rank projection operators, $U_{M}:\mathbb{H}_{r}\to\mathbb{H}_{r}$ where $U_{M}:=\mathcal{D}_{M}\circ\mathcal{E}_{M}$ where $\mathcal{E}_{M}$ and $\mathcal{D}_{M}$ are as defined in section 2. From Lemma 3.3, we have that for any $\tilde{\Lambda}_{\Gamma}\in D_{\Lambda}$:
(19) $\lim\limits_{M\to\infty}\sup\limits_{\tilde{\Lambda}_{\gamma}}\lvert\lvert\Lambda_{\gamma}-U_{M}(\tilde{\Lambda}_{\gamma})\rvert\rvert_{\mathbb{H}_{r}}=0.$
@@
(20) $\forall z_{1},z_{2}\in Z,\quad\lvert\lvert\mathcal{G}^{\dagger}_{\text{ext}}(z_{1})-\mathcal{G}^{\dagger}_{\text{ext}}(z_{2})\rvert\rvert_{L^{\infty}}\leq\omega(\lvert\lvert z_{1}-z_{2}\rvert\rvert_{\mathbb{H}_{r}})$
@@
(21) $\sup\limits_{z\in D_{\Lambda}}\omega(\lvert\lvert U_{M}(z)-z\rvert\rvert_{\mathbb{H}_{r}})\leq\epsilon/3.$
@@
(22) $\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}\lesssim\epsilon/3.$
@@ 3.3 Reconstruction error
Note that the output of the DeepONet lies in the space of $L^{2}$ functions. For any $\tilde{\Lambda}_{\gamma}\in D_{\Lambda}$, for the third term in (3.2) we have, $E_{\mathcal{R}}\leq\sup\limits_{\tilde{\Lambda}\in D_{\Lambda}}{\lvert\lvert(\mathcal{R}\circ\mathcal{P}-\mathbb{I}_{\Omega})(\mathcal{G}^{\dagger}_{\text{ext}}\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}}$, where $\mathbb{I}_{\Omega}$ represents the identity map on the space of $L^{2}(\Omega)$ functions. Let $\tilde{\Lambda}_{\gamma}\in D_{\Lambda}$ and denote $g:=\mathcal{G}^{\dagger}_{\text{ext}}(\tilde{\Lambda}_{\gamma})=\mathcal{G}^{\dagger}(\tilde{\Lambda}_{\gamma})=\Phi^{-1}\circ\gamma$, where $\Phi$ is the link function as described in section 2. Since $\gamma$ is a piecewise analytic function, hence it belongs to some $H^{s}(\Omega)$ for $20<s<1/$ 2. It is easy to see that,
@@
where $\tau_{0}^{\perp}:=\tau_{0}-\Sigma_{k=1}^{P}\langle\tau_{0},\tau_{k}^{*}\rangle\tau_{k}$ is the projection of $\tau_{0}$ onto the orthogonal complement of $\mathrm{span}(\tau_{1},\cdots,\tau_{p})$. Following arguments from [44, Lemma 3.2,3.3, 3.4], we have the following Lemma.
Lemma 3.4.
@@
Finally, the upper bounds shown on the three kinds of error, viz. approximator error, encoder error and reconstructor error, prove Theorem 3.1.